OpenStudy (ashy98):

Given the parent functions f(x) = 5x − 1 and g(x) = 3^x − 9, what is g(x) − f(x)?

OpenStudy (ashy98):

A.) g(x) − f(x) = 3^x − 8 − 5^x B.) g(x) − f(x) = 3^x − 10 − 5^x C.) g(x) − f(x) = −2^x − 8 − 3^x<--------- M answer D.) g(x) − f(x) = −2^x − 10 − 3^x

jimthompson5910 (jim_thompson5910):

You're given these two functions? \[\Large f(x) = 5x-1\] \[\Large g(x) = 3^x - 9\] or is it something else?

OpenStudy (ashy98):

Yes thats it

jimthompson5910 (jim_thompson5910):

wait, it's 5x or 5^x ?

OpenStudy (ashy98):

just 5x

jimthompson5910 (jim_thompson5910):

hmm strange. Your answer choices don't have 5x. They only have 5^x

jimthompson5910 (jim_thompson5910):

This is what I get if it was f(x) = 5^x - 1 \[\Large f(x) = 5^x-1\] \[\Large g(x) = 3^x - 9\] Subtract the functions like this \[\Large g(x) - f(x) = [g(x)] - [f(x)]\] \[\Large g(x) - f(x) = [3^x-9] - [5^x-1]\] \[\Large g(x) - f(x) = 3^x-9 - 5^x+1\] \[\Large g(x) - f(x) = 3^x+(-9+1) - 5^x\] \[\Large g(x) - f(x) = 3^x-8 - 5^x\] The answer is actually choice A. That is assuming f(x) = 5^x - 1 is correct

OpenStudy (ashy98):

yeah your right i have it wrong

OpenStudy (ashy98):

Given the parent functions f(x) = 5x − 1 and g(x) = 3^x − 9, what is g(x) − f(x)? g(x) − f(x) = 3^x − 8 − 5x g(x) − f(x) = 3^x − 10 − 5x g(x) − f(x) = −2x − 8 − 3^x g(x) − f(x) = −2x − 10 − 3^x

OpenStudy (ashy98):

im sorry :/

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (ashy98):

so the answer i think is c

jimthompson5910 (jim_thompson5910):

Is this still the same question? If so, then I show how the answer is A

OpenStudy (ashy98):

yes its still the same question, i just have the x's in the wrong places

jimthompson5910 (jim_thompson5910):

what do you mean? Can you post a screenshot of the problem?

jimthompson5910 (jim_thompson5910):

ok I see now

OpenStudy (ashy98):

it is question 1

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 5x-1\] \[\Large g(x) = 3^x - 9\] Subtract \[\Large g(x) - f(x) = [g(x)] - [f(x)]\] \[\Large g(x) - f(x) = [3^x-9] - [5x-1]\] \[\Large g(x) - f(x) = 3^x-9 - 5x+1\] \[\Large g(x) - f(x) = 3^x+(-9+1) - 5x\] \[\Large g(x) - f(x) = 3^x-8 - 5x\] So the answer is still choice A

OpenStudy (ashy98):

okay i understand it now thank you soo much for taking the time to help me!

jimthompson5910 (jim_thompson5910):

no problem